### v1

Abstract
The rapid evolution of deep learning (DL) models and the ever-increasing size of available datasets have raised the interest of the research community in the always important field of vision-based hand gesture recognition (VHGR), and delivered a wide range of applications, such as sign language understanding and human-computer interaction using cameras. Despite the large volume of research works in the field, a structured and complete survey on VHGR is still missing, leaving researchers to navigate through hundreds of papers in order to find the right combination of data, model, and approach for each task. The current survey aims to fill this gap by presenting a comprehensive overview of this aspect of computer vision. With a systematic research methodology that identifies the state-of-the-art works and a structured presentation of the various methods, datasets, and evaluation metrics, this review aims to constitute a useful guideline for researchers, helping them to choose the right strategy for delving into a certain VHGR task. Starting with the methodology used for study selection, literature retrieval, and the analytical framing, the survey identifies and organizes key VHGR approaches using a taxonomy-based format in various dimensions such as input modality and application domain. The core of the survey provides an in-depth analysis of state-of-the-art techniques across three primary VHGR tasks: static gesture recognition, isolated dynamic gestures and continuous gesture recognition. For each task, the architectural trends and learning strategies are listed. Additionally, the study reviews commonly used datasets —emphasizing on annotation schemes— and evaluates standard performance metrics. It concludes by identifying major challenges in VHGR, including both general computer vision issues and domain-specific obstacles, and outlines promising directions for future research.
keywords:
Vision-Based Hand Gesture Recognition; Human-Robot Interaction; Human-Computer Interaction; Sign Language; Deep Learning; Gesture Recognition Taxonomy
1 Introduction
Understanding human behaviors—such as hand gestures [2], actions [156], and facial expressions [150]—is a key area of artificial intelligence research. There is growing interest in developing Intelligent User Interfaces (IUIs) that offer more natural and intuitive ways for humans to interact with machines, replacing traditional tools like the keyboard and mouse. This has led to increased focus on Hand Gesture Recognition (HGR) across various applications. For example, HGR is used in human-robot interaction in industrial settings (e.g., robotic arm control [134, 96]), healthcare (e.g., robot-assisted surgery [139]), and autonomous mobile robot navigation (e.g., UAVs [194] and UGVs [17, 105]). It also plays a key role in human-computer interaction, especially in AR/VR environments [32, 199, 5]. In addition, the large population of deaf and hard-of-hearing individuals has driven research into automatic sign language recognition systems, aiming to improve communication and inclusion for these communities [196, 26].
Given the importance of the applications mentioned above, Hand Gesture Recognition (HGR) has grown into a broad research field, generally divided into two main approaches: vision-based and sensor-based. Vision-based HGR relies on visual data such as RGB and depth images, as well as features derived from them—like pose heatmaps or optical flow. Sensor-based HGR, on the other hand, uses data from devices such as wearable gloves or surface EMG (sEMG) sensors, and even non-wearable sources like acoustic signals. Researchers have also explored multi-modal methods that combine inputs from different sources. In this study, we focus on vision-based HGR, as sensor-based methods often involve high cost and complex setup, making them less practical for real-world use. Vision-based approaches better align with the goal of HGR: to enable more natural, unobtrusive, and accessible human-machine interaction.
Many of the published surveys consider both sensor-based and vision-based methods together [140, 162, 64, 6]. As a result, the analysis of the proposed methods in these surveys tends to be superficial, because there are so many methods to cover. Other review papers also present the DAQ devices used by researchers [43], further broadening the range of related publications. The undeniable importance of sign language recognition has led some researchers to conduct surveys dedicated exclusively to this topic [2, 155, 170, 145, 174]. In addition, many authors consider both deep learning (DL) and traditional machine learning (ML) methods, or in some cases only traditional ML approaches [174, 140, 43, 132, 64, 6], while others include outdated methods [162, 123, 132, 7, 3, 84]. Although this broader approach helps lay the foundational understanding of VHGR and facilitates a historical perspective, it often limits the focus on current state-of-the-art methods, which are arguably more critical for driving the field forward.
Some surveys aim to provide a quantitative overview of the field [132, 64, 3, 84], offering bibliometric analyses on various aspects of VHGR. While such surveys present a statistical and objective perspective and help identify major subfields, they fall short in advancing the field due to their limited examination of the existing methodologies.
In contrast, to the aforementioned surveys, this survey emphasizes very recent publications and provides a detailed analysis of their contributions, aiming to identify current challenges and innovations that could serve as a valuable starting point for future research. A comparison with existing reviews is presented in Table 1.
Table 1: Comparison of this survey with existing ones. ML refers to traditional Machine Learning/Computer Vision techniques.
Publication Year Range # Papers Input Data Models Scope
[2] 2021 - - Vision-based DL CSLR
[123] 2021 2014-2020 98 Vision-based DL & ML General
[7] 2021 2014-2021 138 Sensor- & Vision-based DL SLR
[3] 2021 2001-2021 649 Vision-based DL & ML SLR
[43] 2022 - - Sensor- & Vision-based DL & ML SLR
[132] 2022 2000-2020 571 Sensor- & Vision-based DL & ML General
[84] 2022 2010-2021 80 Vision-based DL & ML SLR
[155] 2023 2015-2023 - Vision-based DL ISLR
[162] 2024 2014-2024 - Sensor- & Vision-based DL & ML General
[64] 2024 2018-2023 256 Sensor- & Vision-based ML General
[6] 2024 2015-2024 58 Sensor- & Vision-based DL & ML SLR
[174] 2024 - - Sensor- & Vision-based DL & ML SLR
[140] 2024 - - Sensor- & Vision-based DL & ML General
This survey 2025 2021-2024 290 Vision-based DL General
More specifically, the survey focuses on vision-based methods for three key Hand Gesture Recognition (HGR) tasks, which are discussed in detail in the following sections: static HGR [208, 131], isolated dynamic HGR (IDHGR) [82, 68, 214, 60, 206], and continuous dynamic HGR (CHGR) [122, 27, 74]. Static HGR involves recognizing gestures from single frames or spatial data only. Isolated dynamic HGR extends this by analyzing short, pre-segmented sequences to capture both spatial and temporal patterns. Continuous dynamic HGR goes a step further by detecting multiple gestures within long, unsegmented video sequences. Each task presents its own unique challenges. For example, CHGR often relies on weak supervision, making optimization more difficult. At the same time, all tasks share common issues such as interference from irrelevant background movements or variations in lighting, viewpoint, and hand appearance.
Vision-based Hand Gesture Recognition (HGR) has a long history, beginning with traditional computer vision techniques based on hand-crafted features and evolving into complex deep learning pipelines [124]. Early approaches typically followed a three-stage process: pre-processing, feature extraction, and classification. Pre-processing often involved segmenting the hand region using skin color or other visual cues. Feature extraction relied on descriptors like Histogram of Oriented Gradients (HOG), while classification was usually performed using models such as Support Vector Machines (SVMs) or Hidden Markov Models (HMMs), depending on the task. With the rise of deep learning, these hand-crafted features were gradually replaced by learned representations, which offer better generalization. Modern deep learning architectures have become increasingly deep and sophisticated, especially for dynamic HGR tasks. These models first capture short-term motion patterns and then aggregate them to understand long-term context [122, 73, 62, 75, 74]. Due to their superior performance, this study focuses primarily on deep learning-based methods.
Given the broad scope of vision-based Hand Gesture Recognition (HGR) within computer vision, it is crucial to identify the current state-of-the-art methods for the three tasks mentioned above, as well as the key challenges that have shaped the field’s development. This survey aims to 1) provide an overview of the field, enabling readers to quickly find the most relevant publications for their needs, 2) offer a detailed analysis of each method, highlighting their key contributions, and 3) present the state-of-the-art solutions, based on the previous analysis, and explain the main challenges that continue to drive research in HGR.
The main contributions of this study can be summarized as follows:
  1. 1.
    A structured and comprehensive methodology for reviewing the literature;
  2. 2.
    A broad overview of vision-based HGR, organized according to multiple taxonomy criteria;
  3. 3.
    A detailed analysis of recent methods, highlighting their key characteristics and innovations;
  4. 4.
    A synthesis of current state-of-the-art approaches along with a discussion of the key challenges facing the field.
The structure of this paper is organized to guide the reader through a comprehensive review of vision-based hand gesture recognition (VHGR). Section 2 presents the literature review methodology, outlining the systematic approach used to collect and evaluate relevant works. Section 3 introduces a taxonomy of VHGR methods based on diverse criteria, offering a structured classification of existing approaches. Sections 4, 5 and 6 delve into the three main VHGR tasks –—static, isolated dynamic, and continuous gesture recognition—– discussing representative methods and their underlying principles. Special attention is given to sign language recognition and semaphoric gestures in dedicated subsections. Section 7 presents the main datasets used for the evaluation of the presented methods and Section 8 summarizes the evaluation metrics commonly used in the field. Finally, Section 9 outlines key challenges in Human Gesture Recognition and highlights current solutions, while Section 10 concludes the survey by reflecting on recent trends and highlighting open challenges that shape the future of VHGR research.
2 Literature Review Methodology
The review begins with a clear definition of the research methodology, establishing a structured foundation for the comprehensive exploration of the VHGR landscape. This systematic approach is intended to serve as a valuable reference for researchers and practitioners seeking to navigate and contribute to the field. The initial step involved defining the scope of the study, which informed the formulation of inclusion and exclusion criteria used to guide the selection of relevant publications. Subsequently, targeted queries were conducted across major scientific databases, including Scopus and IEEE Xplore, to retrieve works aligned with the established criteria. The final stage entailed a thorough examination of the selected literature, with the objective of distilling the key innovations, contributions, and methodological insights of each publication. These elements are further elaborated in the following sections.
2.1 Survey Scope
Positioned at the intersection of computer vision and human-computer interaction, this work sets out to map the evolving landscape of vision-based hand gesture recognition. By distilling the most impactful recent contributions, it seeks to clarify the field’s current trajectory and highlight its most pressing research needs. The review is structured around five foundational pillars:
  • Advanced deep learning methods introduced in recent literature;
  • Benchmark datasets employed across various VHGR tasks;
  • Established evaluation metrics used to measure system effectiveness;
  • Key challenges that continue to shape the research agenda;
  • Innovative solutions and future directions aimed at overcoming existing limitations.
2.2 Selection Criteria
To ensure the quality, relevance, and completeness of this review, a well-defined set of inclusion and exclusion criteria was established prior to initiating the literature search. The selection process was guided by three key axes:
  • Relevance: Only works that directly address VHGR were considered. This includes methods focused on hand gesture recognition and sign language recognition based solely on visual modalities such as RGB images, infrared (IR), optical flow, motion history, depth maps, and pose representations (e.g., skeleton landmarks and heatmaps derived either from RGB-based pose estimators or visual data acquisition devices). Studies centered on Human Action Recognition (HAR), sensor-based techniques (e.g., using electromyography or radar), or multimodal approaches combining visual and non-visual data were excluded to maintain a focused scope.
  • Up-to-date: To reflect the current state of the field, priority was given to publications from 2021 onward, particularly those introducing new methods. Exceptions were made for seminal works—defined here as milestone contributions post-2015 with high citation counts—that have significantly shaped VHGR research. In the case of datasets, no strict publication date was imposed, but only those actively used by the community or released within the past four years were included. All survey papers considered were published after 2021.
  • Impact: To further focus on seminal and influential works, publications were sourced from highly cited venues (i.e. having an h5-index of 90 or above on Google Scholar), with a single exception for the IEEE International Conference on Automatic Face & Gesture Recognition, which is closely aligned with VHGR. In addition to the above, and aiming to provide the theoretical context of the study, a small number of milestone papers has also been referenced, though not analyzed in the main methods section.
2.3 Literature Search
To identify publications that satisfy the previously defined criteria, a comprehensive literature search was conducted using specialized queries across several well-established scientific databases, including Scopus, IEEE Xplore, ACM Digital Library, and arXiv for pre-prints. Google Scholar was also used to double-check for publications from major venues that might have been overlooked. The search queries were designed based on the previously outlined inclusion/exclusion criteria, utilizing specific commands (e.g., EXACTSRCTITLE, LIMIT-TO) and boolean operators (e.g., AND, OR, NOT) to refine the results. A list of the keywords that were applied in the title, abstract and index terms of each one publication, can be found in Table 2. This process retrieved over 900 publications in the Scopus database. Out of them, only 210 papers were related to vision-based HGR. A second query has been performed on IEEE Xplore, excluding the non-IEEE venues, and resulted in 633 articles, most of them already listed in Scopus. A few simpler queries were performed on Google Scholar and arXiv, and duplicates were ignored, resulting in 290 papers in total, including review papers, methods, and datasets.
Table 2: Keywords used in literature search
Keywords
gesture recognition, gesture classification, sign language recognition, sign language understanding, pointing recognition, human finger pointing, hand gestures, human gestures, gesture, hand gesture, body gesture
2.4 Categorizing the Identified Articles
After applying the inclusion and exclusion criteria, the remaining publications were carefully reviewed for relevance. Those that did not meet the criteria were excluded from further analysis. The selected papers were then categorized into three distinct groups: methods, datasets, and review papers. Each category was analyzed separately, and the corresponding information was organized into three dedicated tables. Publications relevant to more than one category were included in each applicable table as separate entries. Key information for each paper was systematically recorded in the appropriate table columns, as detailed below:
  • Methods: The name of the proposed method, datasets used for evaluation, a brief description of the architecture, the input modality, the learning strategy, and the evaluation metrics;
  • Datasets: For each dataset, we recorded the following details: name, type of gesture based on temporal characteristics (e.g., static, isolated dynamic, or continuous), data modality, total number of samples, number of subjects, and—when applicable to continuous HGR—the number of gesture instances. For sign language understanding (SLU) datasets, we also noted the language, vocabulary sizes (both sign and text), and total duration (for video-based datasets). Additional attributes include the source of the dataset (e.g., laboratory, television, or web), frame rate, resolution, and the environmental conditions under which the data was recorded.
  • Surveys: For survey papers, a concise description of the content was collected.
(a) The distribution of the selected papers on the top-10 venues, in terms of number of publications.
(b) The number of selected publications grouped by year.
Figure 1: Bibliometric statistics for the selected papers.
Additionally, for each publication, metadata such as title, authors’ names, venue, publication year, citation count, keywords, task, application domain, authors’ motivation, challenges, and future directions were also recorded. This meticulous process resulted in a curated collection of 290 papers. Some bibliometrics are provided in Fig. 1. Notably, the number of review papers is limited, highlighting the research gap that this survey aims to address.
3 Taxonomy of Hand Gesture Recognition Methods
Figure 2: Taxonomy of hand gesture recognition methods.
This section performs an overview of the landscape of vision-based hand gesture recognition methods, including various aspects, such as input, representation, applications, etc.. More specifically, a set of complementary and diverse criteria, each focusing on key aspects of a VHGR system, is considered, resulting into the classification of related works into a corresponding set of main categories (while each class can be further divided into sub-categories, as will be detailed below). The different criteria that are used to organize the research works in the field are graphically illustrated in Fig. 2 and are detailed as follows:
  • Application Domain: This criterion refers to the various domains in which VHGR (Visual Hand Gesture Recognition) solutions have been applied. The following categories are considered:
    • Sign Language Understanding (SLU): Involves the automatic interpretation of gestures in either a single sign language (monolingual SLU [102, 12]) or across multiple sign languages (multilingual SLU [126]).
    • Human-Computer Interaction (HCI): Encompasses any form of interaction between users and computers. Numerous studies focus on HCI [131, 193], with particular emphasis on AR/VR applications [199, 32, 5].
    • Human-Robot Interaction (HRI): Includes gesture-based control of mobile robots such as UGVs [17, 152] and UAVs [211, 194], industrial robotic arms [96], assistive robots [29], and surgical robots [139].
    • Human-Machine Interaction (HMI): Focuses on interfaces between humans and machines, especially in automotive contexts. VHGR enables recognition of traffic gestures (e.g., from police officers) [45, 59, 185] and supports the development of more intuitive in-vehicle user interfaces [124, 138, 53, 115, 52].
  • Task: Complementing the application domains, this criterion focuses on the nature of the gestures themselves, specifically their temporal characteristics. Gestures are broadly classified as either static or dynamic, depending on whether they involve a fixed hand posture or a sequence of movements over time. Dynamic gestures typically consist of three temporal phases: preparation, core (or nucleus), and retraction [124], which reflect the gesture’s evolution from initiation to completion. Illustrative examples are provided in Fig. 3. The recognition of dynamic gestures can be further divided into isolated and continuous, considering only one gesture per data sequence and identifying gestures in a continuous manner, respectively.
    Figure 3: Gesture recognition based on task: a) static [33] and b) dynamic [27] VHGR.
  • Input Modality: In continuation of the previous criteria, which focused on application contexts and gesture types, this criterion examines the nature and variety of visual input data used for VHGR. The choice of input modality directly influences the representation architecture and the deep learning models employed. While RGB frames remain the most commonly used modality, other forms of visual data have also been explored—such as depth maps, infrared (IR) images, skeleton joints (and derived features like bones and motion [38]), pose heatmaps [27, 214], optical flow, and motion history [119]. Representative examples are shown in Fig. 4.
    Based on the type of input, methods are broadly categorized as:
    • Pose-based (or skeleton-based): These approaches exclusively process skeleton data.
    • Appearance-based: These rely on visual cues from RGB, depth, IR, or similar modalities.
    (a) From left to right: RGB, Optical Flow, Depth and IR [124]
    (b) Skeleton joints [102]
    (c) Pose heatmap [27]
    (d) Motion history [119]
    Figure 4: Modalities that are considered for VHGR.
    In terms of the number of input modalities, methods can be classified as:
    • Uni-modal: Utilize a single type of input data.
    • Multi-modal: Integrate two or more modalities. These methods can be further categorized based on their multi-modality fusion strategy [177, 60] (see Fig. 5):
        Early (data-level) fusion: Combines modalities at the input level before processing [49, 180].
        Late (score-level) fusion: Fuses final prediction outputs from independent uni-modal models [119, 167, 51].
        Intermediate (feature-level) fusion: Merges features extracted from separate modalities at a single processing level  [172].
        Multi-level fusion: Enables interaction across modalities at multiple stages of processing [214, 60].
    Figure 5: Multi-modality fusion techniques: a) Early (data-level) fusion, b) Late (score-level) fusion, c) Intermediate (feature-level) fusion, and d) Multi-level fusion. Illustration from [60]
  • Type of Information Conveyed: Building upon the previous criterion, which focused on the nature of the input, this dimension considers the meaning or intent behind the gestures. Specifically, it examines the kind of information gestures are meant to transmit [16], resulting in the following categories (illustrated in Fig. 6):
    • Semaphoric gestures convey specific commands or instructions and are commonly used in gesture-based interfaces [124, 131, 17].
    • Language gestures function similarly to semaphoric ones but are exclusively used in the context of sign languages [12, 158].
    • Deictic (pointing) gestures are used to indicate objects, directions, or spatial targets [127].
    • Manipulative gestures aim to interact with or mimic the manipulation of physical objects [50].
    (a) Semaphoric [86]
    (b) Language [12]
    (c) Pointing [127]
    (d) Manipulative [50]
    Figure 6: Main categories of hand gestures based on type of information conveyed.
  • Camera Point of View: Complementing the previous criteria on gesture content and input modalities, this dimension focuses on the spatial relationship between the camera and the performer [199]. The camera’s perspective influences both the visual characteristics of the gesture and the recognition model design. Three primary viewpoints are considered (see Fig. 7):
    • First-person view (egocentric): The camera is mounted on the performer, capturing the gestures from their own perspective.
    • Second-person view: The camera acts as the direct recipient of the gestures, typically simulating the perspective of a human observer or interacting agent.
    • Third-person view: The camera observes the gesture from an external position, detached from the performer.
    (a) First-person view (egocentric) [199]
    (b) Second-person view [149]
    (c) Third-person view [158]
    Figure 7: Different camera points of view for VHGR.
  • Architecture: Building on the previous criteria, which focused on the nature of input and camera perspective, this criterion concerns the architectural design of the gesture recognition system, regarding the modeling stages performed by the backbone. Specifically, it refers to how spatial and temporal information is processed and integrated to model gesture dynamics. Three main categories are identified:
    • Spatial architectures consider only spatial data, such as isolated RGB frames, and fed them into corresponding DNNs (mainly 2DCNNs) to capture spatial information.
    • Temporal architectures extract only temporal features, utilizing sequential models to process sequences of raw data, such as coordinates of skeleton joints.
    • Spatiotemporal modeling architectures, that simultaneously analyze both the spatial and temporal domain, such as 3DCNNs.
    • Spatial + Temporal pipelines, that firstly extract spatial features (e.g., using 2DCNNs) and secondly pass them through a sequential model (e.g., BiLSTM) to perform temporal modeling.
    • Short-term Spatiotemporal + Long-term Temporal pipelines, which aim, given a sequence of data, to capture local context (that is, they extract spatiotemporal features given small parts of the overall sequence) and then fed them into a temporal model to aggregate the partial information and obtain the global vector representation.
  • Learning Strategy: This criterion explores the different training methodologies adopted by researchers to develop VHGR models, often influenced by data availability, task complexity, and desired generalization. The main strategies identified include:
    • Transfer Learning, where models are pretrained on large-scale datasets (e.g., Jester [117]) and subsequently fine-tuned for specific downstream VHGR tasks [112].
    • Weakly Supervised Learning, often used in Continuous Hand Gesture Recognition (CHGR), where temporal annotations are coarse or unaligned, lacking exact gesture boundaries [188].
    • Knowledge Distillation, in which one part of the network (often a teacher model) transfers knowledge to another (student model) to enhance feature extraction and efficiency [122, 197, 27].
    • Contrastive Learning, used either in a CLIP-based framework to support zero-shot/few-shot recognition by leveraging textual descriptions [22, 146, 23], or to improve the discriminative capacity of learned representations [203, 204].
    • Multi-task Learning, where a single model is optimized to perform multiple related tasks simultaneously, encouraging knowledge sharing and improved generalization [102].
    • Reinforcement Learning, applied to optimize decision-making processes, such as keyframe selection, by rewarding effective representations [46].
    • Few/Zero-shot Learning, which enable the model to generalize to new gestures or languages with limited or no annotated examples, often leveraging semantic or textual guidance [22, 146, 23].
    • Adversarial Learning, used as a data augmentation strategy [13] to improve robustness, as demonstrated in adversarial video synthesis networks [128].
    • Masked Reconstruction Pretraining, inspired by BERT-style architectures, in which models are pretrained to reconstruct masked skeleton sequences, injecting structural prior knowledge (e.g., hand topology) [202, 71, 68].
This comprehensive taxonomy of vision-based hand gesture recognition (VHGR) methods, has been grounded in a diverse set of classification criteria that span from application domains to architectural and training strategies. Among these, the task criterion, distinguishing between static and dynamic gestures based on their temporal properties, emerges as particularly pivotal. It fundamentally shapes not only the selection of appropriate models and input modalities but also the interpretation of gestures in real-world scenarios. Due to its central role in determining the complexity and design of recognition systems, the task criterion serves as a natural point of departure for further analysis. The next three sections delve deeper into this dimension by systematically examining the core types of VHGR tasks—Static Visual Hand Gesture Recognition, Isolated Dynamic VHGR, and Continuous VHGR—each with its own challenges, methodologies, and applications.
4 Methods for Static VHGR
Static VHGR aims to identify gestures relying only on spatial data. As this approach neglects temporal information, is only suitable for recognizing static hand gestures. Proposed methods for static VHGR can be further categorized into single-stage and multi-stage approaches, based on whether the architecture consists of a single model or a multi-model processing pipeline, respectively. Indicative works of each category are briefly analyzed into the two following subsections. The third subsection concludes with the main features of state-of-the-art static VHGR methods, and discusses the open challenges that concern the community.
4.1 Single-stage Approaches
Single-stage approaches rely on a single model, which varies depending on the task. For gesture classification, fine-tuned variants of well-known 2D-CNN architectures are employed, such as VGG and DenseNet, or custom lightweight architectures are trained from scratch. When classification is combined with localization of gestures, variants of YOLO are adopted. As far as it concerns the application, static VHGR methods are used in Sign Language Understanding (SLU) and Human Computer Interaction (HCI). However, a few more general-purpose models have also been proposed:
  • For Sign Language Understanding, the authors in [157] developed G-CNN, a shallow 2DCNN to recognize Indian Sign Language (ISL) signs and slightly outperformed various versions of VGG on a self-collected dataset. The authors in [54] proposed the STN-ND-Net, a combination of a Spatial Transformer Network (STN) that improves robustness to spatial transformations, with a 2D-CNN backbone that uses deconvolution operations to reduce the pixel-wise and channel-wise correlations and eliminates the information redundancy of colored images. The success of vision transformer (ViT) motivated authors of SIGNFORMER[93] to adopt it for ISL recognition, while authors in [161] leveraged both skeleton and RGB data in their dual stream feature extractor and an MLP to perform intermediate fusion. A YOLO variant, namely YOLOv4-CSP[8] has also been developed, to simultaneous detect and identify the hand gesture.
  • For gesture-based Human-Computer Interaction the authors in [198] proposed a lightweight 2D-CNN, named LHGR-Net, which combines a Multiscale Structure (MSS) that aims to make the model scale-invariant via atrous convolution with a Lightweight Attention Structure (LAS) based on CBAM. The authors in [80] proposed 2DPSTPP-Net, a 2D-CNN with four stacked convolutional blocks whose features are pooled using max pooling and a novel method, then concatenated to form the final representation, tailored for HCI scenarios. The authors in [61] modified the YOLOv5 to make it more lightweight and developed a model that can be used for gesture-based user interfaces in smart sports stadiums, improving the experience of spectators and athletes.
  • General-purpose models have also been developed. EDenseNet [171] extends DenseNet by adding convolutional layers between dense blocks. SpAtNet [21] is a lightweight 2DCNN with three streams: two MAFF modules for coarse-grained features and one Interleaved Module for high-level representations; their fused features are fed to an MLP for classification. MMFANet [103] performs gesture detection and recognition using a 2DCNN backbone with task-specific regressor and classifier branches.
4.2 Multi-stage Approaches
Figure 8: Architecture of an indicative multi-stage dual-stream approach for static VHGR named FGDSNet [208].
In contrast to single-stage, multi-stage approaches set up a multi-model processing pipeline. Typically, a hand detector or segmentation model is used to remove the background and normalize the spatial dimensions. These methods are primarily applied in Human Robot Interaction (HRI) and HCI scenarios, as detailed below:
  • Human-Robot Interaction, particularly gesture-based mobile robot navigation, has attracted research interest. URGR [17] is a three-stage approach using YOLOv3 for human cropping, HQ-Net for super-resolution to address long-range scenarios, and a classifier for gesture recognition. Similarly, [153] introduced a two-stage method based on depth data: hand segmentation followed by gesture classification using DRCAM, a 2DCNN with stacked Residual Channel Attention Modules (RCAMs) and dense interconnections.
  • Multi-stage static VHGR methods have also been proposed for Human-Computer Interaction. In [9], a pipeline for egocentric HGR and fingertip localization uses YOLOv2 for hand detection, followed by VGG16 for feature extraction and gesture classification. In [34], a lightweight HCI pipeline segments the hand using DeepLabV3+ or Unet, then classifies the cropped image with a 2D-CNN like MobileNet or EfficientNet. [208] introduced FGDSNet, a two-stage approach with dual-stream segmentation (local detail and semantic branches) and multi-level fusion for feature integration. The segmentation mask and original image are fed into a dual-stream classification module, where mid-level fusion combines features for final classification via an MLP.
4.3 State-of-the-art Methods for Static VHGR
The current SOTA static VHGR methods are summarized in 7, based on input modality, architecture, and performance on public datasets, including HGR1 [88], NUS-I [95], NUS-II [136], ASL-FS [137], Jochen-Triesch II [175], and OUHANDS [118] [^1] For datasets with splits for both subject-independent and subject-dependent modes, only the latter is considered.. Since most methods are evaluated on different or custom datasets, a straightforward performance comparison is not possible.
As shown, most SOTA methods rely on RGB data, with depth-based DRCAM [153] lagging behind RGB methods on ASL-FS. Researchers use attention mechanisms [153, 198, 208], multi-stream approaches [21, 208], and segmentation modules for background removal [208, 34, 153] to improve performance. On OUHANDS [118], ASL-FS [137], and NUS-I [95], single-stage methods outperform multi-stage ones, while on NUS-II [136], FGDSNet [208] with segmentation achieves slightly better accuracy than 2DPSTPP-Net [80]. Thus, background removal does not always improve sign recognition, as single-stage methods often surpass multi-stage models.
Table 3: Static HGR methods and their performance on well-known datasets. The highest accuracies are shown with bold. Methods marked with star (*) perform hand segmentation.
Method Modality Architecture Description Dataset Accuracy
SpAtNet [21] RGB Three-stream 2DCNN, where each stream aims to obtain either high-level or coarse-grained features HGR1 93.33%
ASL-FS 99.67%
NUS-II 96.53%
FGDSNet [208]* RGB Dual stream 2DCNN approach and a custom dual-stream segmentation module to obtain the mask HGR1 94.8%
OUHANDS 97.8%
NUS-II 99.8%
[34]* RGB DL-based Hand Segmenation + Classifier HGR1 98.15%
OUHANDS 89.2%
LHGR-Net [198] RGB BaseNet + MSS (adopting the atrous spatial pyramid pooling) + LAS (based on CBAM) HGR1 93.15%
OUHANDS 98.75%
2DPSTPP-Net [80] RGB Convolutional blocks followed by the proposed pooling method, intermediate features are concatenated to obtain the final representation. NUS-I 100%
NUS-II 99.61%
DRCAM [153]* Depth Attention-based 2CNN with residual connections ASL-FS 93.44%
Gesture-CNN [157] RGB 2DCNN consisting of four convolutional layers Jochen-Triesch II 100%
EDenseNet [171] RGB DenseNet with extra convolution layers ASL-FS 99.99%
NUS-II 99.3%
5 Methods for Isolated Dynamic VHGR
Unlike static VHGR approaches, which often fail to capture temporal dynamics and struggle with motion-based gestures, recent advances in Isolated Dynamic Vision-based Hand Gesture Recognition (IDVHGR) have focused on recognizing individual gesture instances from spatiotemporal inputs such as RGB video sequences or skeleton joint trajectories.
When gestures are part of a sign language, the task is referred to as Isolated Sign Language Recognition (ISLR). A notable distinction is that ISLR models frequently leverage both manual features (e.g., hand movements) and non-manual features (e.g., facial expressions, upper body motion), which can significantly enhance recognition accuracy. As a result, many ISLR approaches go beyond hand-centric analysis to incorporate richer contextual cues.
In the following, we present IDVHGR methods grouped by input modality as (a) unimodal appearance-based, (b) pose-based, and (c) multimodal approaches. Then, we conclude with a summary of key innovations and state-of-the-art (SOTA) performance.
5.1 Unimodal Appearance-based Methods
Unimodal appearance-based methods rely on a single input type—such as RGB, optical flow, infrared (IR), motion history, depth, or depth flow—to classify hand gestures. As shown in Fig. 4, these representations resemble color images, which allows for the application of similar computer vision techniques. These methods can be further categorized based on the architecture of the feature extractor, which typically follows one of three main pipelines: (1) spatiotemporal, (2) spatial + temporal, and (3) short-term spatiotemporal + long-term temporal backbones. The following paragraphs explain each of these architectures in more detail.
5.1.1 Spatiotemporal
Spatiotemporal backbones perform spatial and temporal modeling jointly, typically using 3D CNNs or (2+1)D CNNs—either custom-built or adapted from existing models.
In [182], the authors introduced a custom network, (2+1)D-SLR, composed of five stacked (2+1)D convolutional blocks (i.e., a 2D convolution followed by a 1D convolution). This structure lowers computational cost compared to traditional 3D convolutions, making the model more practical for deployment. To enhance temporal modeling, authors in [20] proposed a trainable key frame selection method based on error correction. VGG16 features are used to identify key frames, which are then preprocessed and fed into a custom 3D CNN for classification.
In [186], a dual-stage multi-stream pipeline for CSL recognition was developed. An adapted EfficientDet model—enhanced with cross-level Bi-FPN connections and a dual-channel spatial attention module (DCSAM)—detects both hands. The full image and cropped hand patches are passed through three parallel 3D-ResNet-18 models (following the P3D scheme), with their features fused and classified via an MLP. Authors in [63] introduced TS3C-Net, a modified 3D ConvNeXt that adds an extra 3×3×3 convolutional branch to each block, enhancing spatiotemporal representation. Finally, [173] addressed limitations in fixed-pipeline designs by combining features from both ResC3D and ConvLSTM using the Selective Spatiotemporal (SeST) module. The fused features are passed through fully connected and softmax layers for classification.
5.1.2 Spatial + Temporal
These backbones first extract spatial features from each frame, typically using 2D CNNs or vision transformers, and then model the temporal dynamics with sequential modules such as GRUs or LSTMs. The resulting sequence embeddings are aggregated, and a classifier (e.g., MLP) outputs the final prediction.
In [65], a method tailored for HCI tasks using only depth data was proposed. The method combined Inception-v3 that extracted spatial features with an LSTM for temporal modeling. For ISLR, authors in [178] introduced a pipeline to recognize Indian Sign Language (ISL) words used by deaf farmers. They combined GoogleNet for spatial encoding with a BiLSTM for sequence modeling.
To reduce inference time, a lightweight ISLR model, called TIM-SLR has been proposed in [183]. The model uses a 2D-CNN for per-frame feature extraction, followed by a Time Interaction Module (TIM)—a zero-parameter unit that fuses frame features into a compact representation. Finally, the Swin Transformer presented in [40] is a two-stage transformer-based model that extracts spatial features per frame, while a causally masked transformer processes the sequence, ensuring that only past and present frame features contribute at each step.
5.1.3 Short-Term Spatiotemporal + Long-Term Temporal
The methods in this category follow a two-stage architecture: the first stage captures short-term spatiotemporal patterns from segments of the input sequence, while the second stage aggregates these features to form a final representation of the entire sequence.
EgoFormer [138] is a model for egocentric dynamic gesture recognition optimized for edge devices, which has been introduced in the context of autonomous driving. EgoFormer combines a custom 3D-CNN with a transformer module and an MLP for final classification. For long-range gesture-based UGV control, authors in [18] proposed a pipeline where a lightweight 2D CNN extracts per-frame features, which are used to select key frames. After cropping the human region, the selected frames are passed through a SlowFast network, followed by a transformer and an MLP.
To address data scarcity in sign language understanding, [22] proposed a zero-shot learning framework. Sign videos are represented using BERT-encoded textual descriptions combined with hand-crafted attributes. This representation guides a dual-stream backbone composed of short-term spatiotemporal modules (e.g., I3D) and long-term models (e.g., BiLSTM).
Figure 9: The zero-shot framework proposed by authors in [22]. Each branch of the dual-stream backbone consists of a spatial-temporal model, such as the I3D, followed by a long-term recurrent model. The BERT is utilized to map textual description into the embedding space. These features are also combined with hand-crafted attributes.
5.1.4 Other
Some appearance-based methods fall outside the previously discussed categories but offer valuable contributions, especially in Sign Language Recogition (SLR). For example, the method proposed in [113] for Indian SLR (ISLR) compresses an entire frame sequence into a single image, which is then processed by a MobileNet. A second MobileNet extracts per-frame features, which are passed through a two-layer LSTM. The outputs of both streams are concatenated and refined using a custom shallow 2D-CNN. A different approach was presented in [94], where the authors introduced a dual-stream ISLR architecture combining Inception V4 with a novel convolutional module called stem.
In the context of pose-based ISLR, [110] proposed a dual-stream model with keyframe selection. A 2D-CNN and attention mechanism filter out less informative frames. The remaining keyframes are processed by two branches: the Hand stream, which crops and analyzes hand regions with attention blocks, and the Full stream, which processes the entire frame. The fused features are used for final classification. To improve signer-independence and reduce overfitting, the SwC GR-MMixer [109], another lightweight model based on MLP-Mixer, incorporates GRU units for temporal modeling and a shifted window concatenation strategy to enhance sequence representation.
5.2 Pose-based Methods
Pose-based methods, also known as skeleton-based approaches, rely on structured motion data—such as skeleton graphs and pose heatmaps—to classify gestures. These methods often incorporate various forms of pose information, including joint coordinates, bones, motion trajectories, and joint angles, sometimes in combination [38]. They also explore joint subsets related to specific body regions [135], often adopting multi-stream architectures to handle these diverse inputs.
Given the natural graph structure of skeletal data, many methods employ Graph Convolutional Networks (GCNs) originally developed for human activity recognition. In the sections that follow, we group the approaches based on the structure of their backbone, following the same organization as in the previous subsection. Notably, pose information alone is often insufficient for accurate sign recognition [125]. Consequently, many models integrate additional cues—typically from RGB inputs—using 3D CNNs and perform score-level fusion to enhance performance.
5.2.1 Spatiotemporal
Pose-based methods that incorporate spatiotemporal representations simultaneously process spatial and temporal data to improve sign language understanding (SLU), sign language recognition (SLR), and human-computer interaction (HCI).
In Sign Language Understanding (SLU), pose-based methods aim to capture dynamic hand and body movements for gesture recognition. SKIM [104] uses a dynamic graph convolutional network (GCN) combined with temporal convolution (TCN) to analyze both global and local body regions, with a feature extraction module focusing on high-motion joints. Similarly, MC-LSTM [133] employs multiple streams (face, hand, and whole body) processed by temporal convolutions and GCN for improved feature ensembling. Another notable method, P3D [98], uses a transformer-based model with part-wise and whole-body encoders to process landmark data for classification. For improved performance, HWGAT [135] utilizes hierarchical spatial windows with attention mechanisms to highlight key body parts. TMS-Net [38] proposes a six-stream GCN-based approach for processing skeleton joints, bones, and motion data, with inter-stream information exchange. Additionally, methods like SLGTformer [168] and VSCT [25] address the challenge of data scarcity by leveraging transformers and few-shot learning, respectively.
In Sign Language Recognition (SLR), pose-based methods often focus on improving the accuracy and robustness of sign classification. Authors in [161] present a dual-stream GCN model for Korean SLR, analyzing joint and motion data. Similarly, MULTISTAGE-CGAR [121] proposes a dual-stream architecture with separable temporal convolution and attention mechanisms for improved feature extraction. For multi-stream fusion, SwC GR-MMixer [109] combines GCN with GRU units for temporal modeling and attention mechanisms to reduce overfitting. Notably, L-TSSI-DenseNet [97] explores skeleton graph representations by converting joint coordinates into image-like matrices processed by DenseNet for classification. Moreover, STST [201] leverages a two-stage transformer-based approach for global and local spatial-temporal feature extraction, offering a dynamic alternative to traditional methods.
In Human-Computer Interaction (HCI), pose-based methods address challenges such as data overfitting and the need for self-supervised learning frameworks. MAE-based frameworks [79] use a masked autoencoder to reconstruct skeleton graphs, focusing on motion and spatial relationships. 3sISTGCN [167] proposes a three-stream model to process joints, bones, and motions, applying modified ST-GCN layers and late fusion for improved classification. Another approach, DG-STA [5], tackles class-incremental learning in HCI by training models on non-overlapping tasks without access to old training data. DSTSA-GCN [30] addresses temporal expression limitations by stacking three stages of graph convolutions and multiscale temporal convolutions. Additionally, DSTA-SLR [77] improves spatiotemporal modeling by introducing dynamic adjacency matrices and multi-stream fusion, enhancing the accuracy of joint, bone, and motion-based recognition tasks.
5.2.2 Spatial + Temporal
Many models use a two-step approach, where the spatial domain is first processed by a spatial perception module to extract spatial features. These features are then passed into a sequential model for temporal aggregation, resulting in a combined spatial-temporal representation of the input sequence. For instance, GCN-BERT [176] utilizes a two-layer GCN for spatial modeling, followed by BERT for temporal modeling.
Other general-purpose methods adopt a multi-stream, two-stage pose-based approach. In [129], three streams, each containing three 1D convolutional layers, are combined with a parallel module of average pooling layers. The extracted features are then concatenated and input into an LSTM, followed by a fully connected layer and softmax for classification. Similarly, STr-GCN [165] employs a two-stage model where a GCN spatial feature extractor is first applied to obtain intra-frame features and then passes the features through a transformer encoder for modeling inter-frame temporal relationships. The final classification scores are generated by an MLP.
5.2.3 Short-Term Spatiotemporal + Long-Term Temporal
In this approach, the backbone first extracts features from smaller segments of the skeleton sequence to capture short-term spatial-temporal relationships. Then, long-term temporal modeling is applied to aggregate global context across the entire sequence.
For example, gesture recognition in autonomous driving has become an important research area. In [59], the authors propose a pose-based approach called MG-GCT for recognizing police traffic gestures. The model uses two graph convolution streams: one to process skeleton joints and another to analyze joint motions. The features extracted from both streams are fused and passed through a transformer encoder for temporal modeling, with the final classification performed by an MLP. In another approach, TD-GCN [107] introduces a pose-based method for IDHGR. Unlike traditional methods that use fixed skeleton topologies, TD-GCN employs temporal-sensitive topology learning to enhance feature quality. First, graph convolution is used to extract high-level spatiotemporal representations, which are then used to calculate temporal and channel-independent adjacency matrices. These matrices help model the skeleton, and the features are fused using the proposed TCF module.
5.2.4 Temporal
Some pose-based methods bypass spatial modeling and focus directly on sequential processing. Typically, keypoints are slightly preprocessed—often by extracting geometric features such as bone angles—and then passed into sequential models like RNNs or TCNs to produce a vector representation of the input sequence.
In [139], the authors proposed a pose-based method for surgical robot teleoperation. Preprocessed skeleton joints are fed into an LSTM, and the resulting features are passed to an MLP for final classification. The model was evaluated on a custom dataset. Similarly, in [151], a cross-domain gesture recognition approach was developed using geometric features (e.g., bone angles), which are processed by a TCN. The resulting features are summarized into a sequence descriptor, and a linear classifier predicts the final class.
To address data scarcity, the authors in [128] introduced AVSN, an adversarial learning-based data augmentation framework. The method jointly optimizes augmentation parameters and the HGR classifier. A BiLSTM serves as the discriminator, while an MLP acts as the generator. The generator produces challenging augmented data that exploits weaknesses in the discriminator, which in turn learns to correctly classify gestures.
5.2.5 Other
Some architectures adopt more complex designs or alternative skeleton representations, placing them outside the previously defined categories.
In [1], the authors combined a Fourier CNN with a BiLSTM to extract spatiotemporal features from 3D skeleton joints and perform ISLR. A trainable binary mask was also used for feature selection to eliminate redundancy. Similarly, authors in [10] performed ISLR using a feature thresholding technique that transforms skeleton features based on their variance and importance. These transformed features undergo exhaustive search for selection and are then processed by a deep CNN for classification.
To address limitations of color-coded feature maps, the authors in [89] proposed the Joint Motion Affinity Map (JMAM) as a new skeleton representation for ISLR. JMAM is input to a custom multi-stream attention-based 2DCNN, which extracts features and performs classification. Key frame selection was explored in [46] through the Frame Selection Network (FS-Net), trained using reinforcement learning. The agent selects frames for classification by DD-Net and receives a reward of +1 or -1 based on the correctness of the classification outcome. Finally, [101] proposed a lightweight dual-stream method to fully capture spatiotemporal features. One stream employs stacked attention-based modules, while the other uses an autoencoder. The outputs of both are classified independently, and their predictions are averaged to yield the final decision.
5.3 Multimodal Methods
Multimodal approaches combine different input types (e.g., RGB, Depth, Skeleton, Text) to improve gesture recognition performance by leveraging complementary information. IDVHGR multimodal methods do not seem to follow any trend, maybe due to the wide spectrum of alternative options (combinations of streams, fusion techniques, modalities, etc.), so we group them by application domain.
5.3.1 General Purpose
In [100], a dual-stream 3DCNN followed by average fusion was proposed for RGB-D gesture recognition. To filter out redundant visual features ($\upsilon$) and retain only the most discriminative components ($z$), a memory bank mechanism was introduced. The training objective maximizes mutual information between $z$ and the gesture class $y$, while minimizing that between $z$ and $\upsilon$. The memory bank is updated via read/write operations during training. Similarly, addressing redundancy, authors in [206] applied Neural Architecture Search (NAS) and attention mechanisms. A modified I3D backbone replaces some inception modules with learnable NAS cells. These progressively transition from small-kernel to dilated convolutions. Dynamic Static Attention (DSA) modules focus on motion and hand/arm regions, aided by a parallel HeatmapNet trained jointly through an auxiliary loss. For RGB-D input, late fusion is used.
To move beyond the limitations of late fusion and exploit modality interactions more effectively, [207] proposed a decoupling-recoupling framework. Independent spatial (with a Focused Region Proposal (FRP) module) and temporal modeling (via attention) are first performed to extract modality-specific features. These are then fused mid-way using a shared MLP-AE module. Multi-level fusion was further studied in [60], where 3DCNN backbones are combined with a Convolutional Transformer Fusion Block (CTFB). Injected at multiple layers, CTFB fuses intermediate features across streams for better spatial-temporal integration.
To address data scarcity, several works explored zero-shot learning (ZSL) by incorporating textual gesture descriptions encoded by pretrained language models (mainly BERT). In [146], a dual-stream model combined a ViT for spatial modeling and an LSTM-MLP for temporal modeling, learning visual features aligned with textual embeddings. Similarly, [147] fused skeleton-derived geometric features with RGB-C3D-LSTM-AE representations into a shared semantic space for classification. To reduce computational cost, [53] replaced the transformer’s self-attention mechanism with convolutional layers and a gated feedforward network, resulting in the lightweight ConvMixFormer architecture.
In [180], knowledge distillation and contrastive learning were employed to enhance multimodal performance. Three 3DCNNs (for RGB, Depth, and Optical Flow) extract features fused via Coordinate Attention (CA). I3D serves as the teacher, and ResNet as the student. CLIP is also integrated for zero-shot gesture representation. The limitations of standard 3DCNNs were explored in [114], where 3D Central Difference Convolution (3D-CDC) replaces standard 3D convolutions. Residual hybrid blocks and a transformer module handle temporal modeling. [195] also adopted 3D-CDC in a multi-stream, multi-rate setup with multi-level fusion. NAS optimizes the configuration of CDC types, inter-stream connections, and frame rates.
5.3.2 Sign Language Understanding
Several recent works have proposed multimodal and pose-centric architectures for Isolated Sign Language Recognition (ISLR), focusing on combining RGB and skeleton data.
A prominent line of research is the SignBERT framework [68], which uses a spatial GCN with temporal and hand chirality encoding to extract gesture features. These are fed into a transformer encoder, while a hand-model-aware decoder reconstructs masked inputs during self-supervised pretraining. Predictions from a parallel RGB network are fused at inference time. An enhanced version, SignBERT+ [71], introduces spatial-temporal position encoding and an improved fusion module for RGB and pose data. The architecture is illustrated in Fig. 10. Similarly, MASA [203] adopts the same GCN+Transformer backbone, trained with contrastive learning on pose reconstruction. BEST [202] extends this by embedding pose triplets (left hand, right hand, body) with BERT and applying late fusion with RGB data.
Figure 10: The pose-based stream of the SignBERT+ [71], an extension of the SignBERT [68], during SSL pretraining. A GCN is first utilized to extract the gesture features, which are after combined with spatial-temporal positing encoding and subsequently pass through a transformer encoder. During pretraining, a MANO-based decoder is placed after the transformer to perform reconstruction of the masked input. The decoder is replaced with a prediction head during downstream fine-tuning.
To improve feature invariance, a contrastive approach was proposed in [204], enforcing consistency among hand, trunk, and motion features. The same GCN+Transformer architecture is used, with late fusion of pose and RGB predictions. The SAM-SLR framework [82] combines six streams (RGB, depth, skeleton) with late fusion of predictions. SAM-SLR-v2 [81] adds a 3D hand graph stream and learnable fusion weights (Fig. 11).
Figure 11: The architecture of SAM-SLR-v2 [81], which is based on the SAM-SLR [82]. The architecture consists of seven streams, regarding various modalities, including RGB frames, optical flow, pose heatmaps and skeleton joints. The predictions of each branch are fused in a learnable late manner.
In VTN-HC/PF [36], hand-cropped images are processed via a 2D-CNN for spatial modeling, followed by a transformer and MLP for temporal and final predictions. The PF variant incorporates geometrical features from pose data. Motion History Images (MHIs) were employed in [119], where a 3DCNN with attention and residual connections analyzes MHIs, while a dual-stream version adds a 2DCNN for RGB data, with predictions fused via weighted summation.
To address data scarcity, ZSSLR [216] adopts a CLIP-based zero-shot framework using RGB and pose data. In [24], a few-shot cross-lingual method is proposed, combining two 3DCNNs and a pose estimator. Features are merged through a Temporal Feature Aggregation Module with self-attention and 1D convolution.
NLA-SLR [214] targets semantically similar signs using a multistream backbone (RGB and pose heatmaps via S3D), followed by a head with language-aware label smoothing and an auxiliary inter-modality mix-up classifier. MSNN [116] uses five I3D streams (RGB and optical flow) plus ST-GCN for pose. hDNN-SLR [142] introduces a hybrid deep network with a 3DCNN for spatial detection, an attention-based BiLSTM for temporal modeling, and a VAE-based module for feature discrimination. Finally, Uni-Sign [102] proposes a unified multimodal architecture with three pose encoders and an RGB encoder fused via an attention-based module. Features are passed through temporal encoders and a pretrained LLM. Training involves pose-only and joint RGB-pose pretraining followed by downstream fine-tuning on ISLR and CSLR tasks.
5.3.3 Human-Computer Interaction
Motivated by the difficulty of recognizing fine-grained gestures in HCI applications, authors in [200] proposed HandSense, a method relying on depth and RGB data. Two 3DCNNs are deployed to extract features for each one modality. These features are fused and fed into a SVM to obtain the final prediction. Authors in [87] proposed an approach relying on RGBD Dynamic Images (RDDIs). These images, are obtained by the combination of RGB and estimated depth data. The resulted RDDI is fed into a modified Xception, resulting from removing the top 4 Xception modules, in order to extract meaningful features. These features are finally used to perform the classification.
5.3.4 Other
Inspired by the success of transformers, GestFormer [51] introduces a lightweight transformer-based framework for gesture-driven HRI. A 2DCNN extracts per-frame features, processed by PoolFormer layers (replacing attention) and two custom blocks for multiscale pooling/WT and token mixing. Each modality is processed independently, and outputs are averaged. In [49], a hybrid 3DCNN/LSTM model fuses RGB, depth, and skeleton data (from an enhanced pose estimator) via weighted sum. The fused input feeds a spatiotemporal model combining 2D/3D convolutions and ConvLSTM layers.
5.4 State-of-the-art Methods for Isolated Dynamic HGR
Recent IDHGR approaches demonstrate high architectural diversity. Pose-based methods often rely on spatial GCNs and temporal transformers [68, 71, 202, 204, 203, 165], or on ST-GCN variants [107, 79, 167]. Due to their limited standalone performance—mainly caused by pose estimation failures in challenging conditions (e.g., motion blur) [203]—pose-based models are frequently fused with simple RGB-based branches (e.g., I3D), using late fusion via score averaging [68, 71, 202, 204, 203, 104].
Appearance-based methods typically follow a multi-stream 3DCNN design to process individual modalities (e.g., RGB, depth, optical flow) [180, 206, 60, 100, 119, 195], employing late or multi-level fusion. Notably, SAM-SLR [82] and SAM-SLRv2 [81] show that adding more modalities—even with simple backbones—can significantly improve accuracy.
Advanced learning strategies further improve performance. Self-supervised learning (SSL) via masked pose reconstruction [202, 68, 71, 79, 203] enhances representation quality by injecting hand topology priors. Neural Architecture Search (NAS) has been used to discover optimal network structures [206, 195], while multi-task learning, as in Uni-Sign [102], boosts performance by leveraging auxiliary tasks and larger training corpora. Performance comparisons across benchmarks are summarized in Tables 4, 5, and 6.
Table 4: Brief description and reported accuracy of state-of-the-art methods on IsoGD [179].
Method Modality Architecture Description Training Setting Accuracy
3SCKI [180] RGBD, OF Three-stream 3DCNN with mid-level fusion Knowledge distillation between two instances of the pipeline, CLIP 65.87%
SlowFast+NAS [195] RGBD, OF Multi-stream and multi-rate 3DCNN with 3D-CDC and multi-level fusion NAS 66.23%
RAAR3DNet [206] RGBD Dual stream I3D with adoptable cells, spatial and temporal attention Multi-task loss and NAS 66.62%
De+Recouple [207] RGBD Dual-stream spatial + temporal with AE for multi-level fusion and mechanisms to reduce redundancy Self-distillation for inter-frame learning, extra supervision to enhance the AE 66.79%
CTFB [60] RGBD Dual-stream I3D with multi-level fusion - 69.02%
[100] RGBD Dual-stream Res3D-18 + average fusion Memory network to enhance the discriminative capability of the extracted features and information-theoretic loss function 74.08%
Table 5: Performance of uni-modal pose-based methods on SHREC [166] and DHG [37].
Method Architecture Split SHREC DHG
[120] Three-stream approach: attention-based spatial-temporal model, attention-based temporal-spatial model and MLP 14 97.01% 92.00%
28 92.78% 88.78%
MS-ISTGCN [167] Three-stream model: ST-GCN with adoptive spatial GC, extended temporal GC and attention 14 96.7% 93.7%
28 94.9% 91.2%
TD-GCN [107] ST-GCN variant with channel-dependent and temporal-dependent adjacency matrices 14 97.02% 93.9%
28 95.36% 91.4%
[101] Two-stream approach consisting of an attention-based dual branch module and MAE 14 96.9% 94.21%
28 94.17% 92.11%
SBI-DHGR [129] Multi-stream 1DCNN + BiLSTM 14 97% 94.64%
28 92.36% 91.79%
STST [201] Two-stage transformer-based approach 14 97.62% 94.82%
28 95.83% 93.18%
DSTSA-GCN [30] Three stacked stages, each one consisting of group channel-wise and temporal-wise GCN, followed by multiscale TCN 14 97.74% 95.04%
28 95.37% 93.57%
DRL-DDNet [46] DD-Net for classification and FS-Net for keyframe extraction 14 96.5% -
28 93.8% -
Table 6: Performance of the current SOTA methods on ISLR. PI, PC, OF and DF stands for per-instance accuracy, per-class accuracy, optical flow and depth flow, respectively. Methods marked with star (*) use additional training data.
Method Modality Architecture Description Training Setting Dataset PI PC
BEST [202] RGB, Pose d-VAE + GCN + Transformer & RGB-based encoder + late fusion SSL pretraining (masked tokens reconstruction) MSASL 71.21% 68.24%
WLASL 54.59% 52.12%
CSL500 97.7% -
SignBERT [68] RGB, Pose Spatial GCN, temporal and chilarity embedding + Transformer & RGB-based stream + late fusion SSL pretraining (MANO-based masked hand reconstruction) MSASL 71.24% 67.96%
WLASL 54.69% 52.08%
CSL500 97.6% -
SignBERT+ [71] RGB, Pose Extension of SignBERT with Spatial-Temporal Position Encoding and enhanced late fusion technique The same as SignBERT MSASL 73.71% 70.77%
WLASL 55.59% 53.33%
CSL500 97.8% -
MASA [203] RGB, Pose Spatial GCN + Transformer & RGB-based stream SSL pretraining (masked input reconstruction), contrastive learning WLASL 55.77% 53.13%
TMS-Net [38] Pose Multi-stream decoupled GCN with multi-level fusion - WLASL 56.4% -
SKIM [104] Pose Multi-stream channel-wise multi-scale GCN + late fusion Hand part mixing for data augmentation WLASL 57.73% 55.40%
[204] RGB, Pose GCN+Transformer and I3D + late fusion Contrastive learning to enhance invariance, consistency constraints between features of hand and trunk parts and between joints and motions MSASL 74.18% 72.18%
WLASL 58.06% 55.66%
CSL500 97.8% -
SAM-SLR-v2 [81] RGBD, Pose, HHA, OF, DF Extension of SAM-SLR-v2 with 3D graph stream and learnable weights for late fusion - WLASL 59.39% 56.63%
CSL500 99.00% -
NLA-SLR [214] RGB, Pose Four-stream network based on S3D with two frame rates and multi-level fusion Language-aware label Smoothing and inter-modality mixup with an auxiliary classifier MSASL 73.80% 70.95%
WLASL 61.26% 58.31%
Uni-Sign [102]* RGB, Pose A pretrained LLM process features extracted by pose, visual and temporal encoders Three-stage training, multi-task learning with additional training data MSASL 78.16% 76.97%
WLASL 63.52% 61.32%
6 Methods for Continuous Dynamic HGR
In contrast to isolated dynamic hand gesture recognition (IDHGR), where each sample contains at most one gesture, continuous hand gesture recognition (CHGR) deals with unsegmented gesture sequences. The goal is to recognize gestures from untrimmed videos and generate the corresponding gesture sequence. CHGR is typically treated as a weakly supervised task, as the precise temporal boundaries of gestures are not provided; the model learns to identify them implicitly, often using Connectionist Temporal Classification (CTC) loss [188, 55].
While CHGR has been explored in domains like autonomous driving [45, 185], most research focuses on Sign Language Understanding (SLU), where it is commonly referred to as Continuous Sign Language Recognition (CSLR) or sign2gloss. CSLR aims to produce a sequence of glosses [^2] Glosses are atomic lexical units in sign language, analogous to words in spoken language. from video input. Due to the importance of non-manual features (e.g., facial expressions and upper-body movements), many approaches also incorporate these regions to improve recognition accuracy. Some CSLR methods are further extended to support gloss-based sign language translation (SLT) by adding a gloss-to-text module [56, 215, 27].
In the following, we review recent advances in CHGR, grouped by input modality. We highlight key architectures, learning strategies, and benchmark results.
6.1 RGB-based
Most state-of-the-art RGB-based Continuous Hand Gesture Recognition (CHGR) models follow a three-stage pipeline architecture: (1) a spatial model (commonly a 2DCNN or a transformer) to extract frame-wise features; (2) a short-term temporal model (mainly 1DCNN or TCN) for local temporal modeling and gloss-level feature extraction; and (3) a contextual module (e.g., BiLSTM) for long-term sequence modeling and final gesture sequence prediction [73, 122, 62, 74, 197, 70, 75, 72, 76]. This standard architecture was introduced in [31]. This decoupled spatial-temporal structure has been favored over unified spatiotemporal models like 3DCNNs due to the latter’s inability to define clear temporal boundaries [2].
Across the various methods that extend the initial three-stage pipeline, six broad enhancement strategies emerge: 1) Auxiliary Supervision & Multi‐Task Objectives, which enrich the standard three‐stage pipeline with extra losses and segmentation tasks; 2) Specialized Temporal Modeling, replacing or augmenting the 1D‐CNN with learned pooling, multi-scale fusion, or hierarchical boundary detectors; 3) Attention & Fusion Modules, inserting correlation, self‐emphasis, or transformer‐fusion blocks into the visual encoder; 4) Backbone Replacements & Hybrids, swapping 2D-CNNs for vision transformers or hybrid CNN–attention architectures; 5) Incorporating Linguistic Priors, aligning visual features with gloss embeddings via cross-modal decoders or contrastive losses; and 6) Data Scarcity & Multilingual Strategies, leveraging self-supervision, prototype memories, and automatic dictionary construction to overcome limited annotations and support multiple sign languages.
Many RGB‐based CHGR methods build on the canonical 2DCNN + 1DCNN + BiLSTM architecture by introducing auxiliary losses and multi‐task objectives. For example, VAC‐CSLR [122] augments the visual and temporal modules with a Visual Enhancement (VE) loss and a Visual Alignment (VA) distillation loss to accelerate convergence and improve representation quality (see Fig. 12). Similarly, SMKD [62] adds two auxiliary classifiers—one after the 1DCNN and one after the BiLSTM—to perform synchronous gloss segmentation alongside recognition, using a three‐stage training schedule to mitigate gradient “spikes.” C2SLR [213] enforces Spatial Attention Consistency (SAC) by focusing the visual module on pose‐informed regions, and Sentence Embedding Consistency (SEC) to align the visual and sequential representations of the same gloss.
Figure 12: An overview of the VAC-CSLR [122]. The feature extractor provides local information, that is processed by the subsequent alignment module to predict the gloss sequence. $F_{a}$ and $F_{b}$ denote auxiliary classifiers, while $L_{VE}$, $L_{VA}$ and $L_{CTC}$ the VE, VA and CTC loss, respectively. This training setting constitutes the baseline for most of the proposed methods.
To better capture local temporal patterns, several works replace or enhance the standard 1D-CNN. Temporal Lift Pooling (TLP) [73] adopts a learned pooling scheme inspired by the lifting transform, eliminating handcrafted kernels. mLTSF‐Net [192] couples content‐aware feature selection with position‐aware convolutions to fuse adjacent‐frame similarities at multiple scales. HS‐I3D [190] extends I3D by extracting intermediate embeddings and deploying a U‐Net–style decoder that upsamples temporally, enabling simultaneous sign detection and boundary localization.
Other methods inject lightweight attention and correlation blocks into the spatial encoder. CorrNet [74] adds parallel correlation and identification modules after each 2DCNN stage to model inter‐frame trajectories and highlight informative regions; CorrNet+ [76] further replaces its correlation operator with a more efficient variant and inserts an attention layer. SEN‐CSLR [75] introduces self‐emphasizing spatial (SSEN) and temporal (TSEN) modules to reweight feature maps and keyframes with minimal overhead. Convolutional Transformer Fusion Blocks (CTFB) [60] fuse intermediate features across 3DCNN branches to enhance cross‐modal integration.
A growing number of approaches swap the 2D-CNN for transformer‐based encoders or hybrid CNN–attention modules. CLIP‐SLA variants [14] leverage a frozen CLIP visual encoder, augmented with LoRA or MLP adapters and an auxiliary classifier for early supervision. Swin‐MSTP [15] pairs a Swin transformer with an MS‐TCN that employs TLP and dilated convolutions, followed by a BiLSTM for global context. TCNet [111] enhances a CNN backbone with trajectory and correlation modules, while SlowFastSign [4] modifies Slow‐Fast‐101 with bidirectional fusion pathways and auxiliary temporal classifiers to refine feature exchange. Most CSLR methods still rely on CNN backbones and thus struggle to model correlations across spatial regions. To address this, SignGraph [^3] Be aware of an IDHGR method with the same name. [47] treats each frame’s patches as nodes in a dynamic graph and alternates between Local Sign Graph (LSG) and Temporal Sign Graph (TSG) modules to capture intra‐frame and inter‐frame relationships; a multiscale variant, MultiSignGraph, further enhances spatial–temporal context (Fig. 13).
Figure 13: An illustration of the SignGraph [47]. The RGB frames are split into patches. Dynamic graphs are utilized to represent the relations between intra-frame and inter-frame patches, that constitute their nodes.
To exploit gloss co-occurrence and language structure, some models fuse linguistic embeddings with visual features. SignBERT‐HKSL [212] combines a (3+2+1)D ResNet with BERT and BiLSTM, pretraining on masked inputs and iteratively fine‐tuning with masked outputs. C2ST [197] replaces the CTC decoder with a Conditional Gloss Decoder (CGD) that models gloss dependencies, and aligns visual and gloss embeddings via cross‐modal match loss. GPGN [58] introduces a Cross‐Modal Match Loss between BERT‐derived gloss features and PDC‐TFE–extracted visual embeddings, followed by a BiGRU for sequence modeling (see Fig. 14). CVT‐SLR [205] pretrains an asymmetric VAE on a pseudo gloss‐to‐gloss task, then bridges the VAE and visual encoder with a video–gloss adapter and contrastive alignment loss (see Fig. 15). Fully transformer‐based PiSLTRc [191] applies content‐aware neighborhood gathering and position‐informed attention for end‐to‐end CSLR. Authors in [72] also introduce a prior-aware cross-modality augmentation that trains a transformer to reconstruct masked gloss sequences and then uses it to generate pseudo video–text pairs by guiding gloss substitutions and insertions.
Figure 14: Authors in GPGN [58] leveraged the linguistic prior from the pretrained BERT. The Cross-Modal Match Loss is introduced, in order to maximize the correlation between the gloss and the visual embeddings.
Figure 15: The approach proposed by authors in CVT-SLR [205]. In the first training stage, a VAE is pretrained on the gloss2gloss pseudo-translation task in order to gain textual knowledge. In the second stage, an adopter is placed between the VAE and the visual encoder, while the consistency of the visual and temporal encoder is maintained via contrastive learning.
Finally, data augmentation and cross‐lingual strategies help alleviate limited annotations. Fangyun Wei et al. [188] build a “dictionary” by using a pretrained CHGR model to automatically segment continuous videos into isolated signs, then mapping German and Chinese sign corpora into a shared embedding space. CSGC [143] maintains a Gloss Prototype memory that stores class prototypes updated during training; a gloss contrastive loss then enforces cross‐sentence consistency.
6.2 Pose-based
Pose-based (or skeleton-based) approaches for CHGR rely exclusively on body pose information, typically in the form of 2D/3D skeleton joints or pose heatmaps. These methods aim to model human motion while discarding appearance-based noise, making them attractive for real-time and privacy-preserving applications.
Several works exploit structured keypoint information by dividing the body into semantic regions such as the face, hands, and upper body. MSKA-SLR [56] processes each region independently via attention-enhanced temporal convolutions across four branches. The resulting features are fused to decode gloss sequences using a CTC loss, complemented by frame-level self-distillation. Similarly, CoSign [83] employs group-specific GCNs for each subset and introduces regularization to preserve inter-stream consistency. Their dual-stream variant integrates joint positions and motion features using multi-level fusion to capture both static and dynamic cues.
Temporal localization of gestures has also been explored in the literature. TY-Net [39] draws inspiration from YOLO to detect gesture instances in time by predicting their boundaries and centers, analogous to spatial bounding boxes.
Beyond joint coordinates, pose heatmaps serve as an alternative input modality. The multi-cue framework in [210] combines spatial and temporal reasoning by localizing facial and hand regions via a pose estimator and subsequently encoding them with a temporal module, before decoding gloss sequences with a CTC head. The FA-STGCN model [45] further improves robustness to motion scale and frequency variability. It normalizes skeleton inputs and applies alternating spatial and temporal GCNs. Spatial modules learn latent joint relationships, while temporal processing is split into global and local paths—extracting coarse-grained and fine-grained motion patterns, respectively.
Finally, in the context of autonomous driving, Wang et al. [185] developed a two-stage pose-based system to classify road intersection gestures. The first stage extracts geometric features and encodes them via LSTM layers. The second stage uses a dual-stream setup, where one branch processes geometric descriptors and the other uses a ConvLSTM to handle co-occurrence matrices. Features are fused with directional orientation indicators for final classification.
6.3 Multimodal
Multimodal approaches for CHGR combine multiple input types—such as RGB frames, pose heatmaps, depth maps, and optical flow—to mitigate the limitations of unimodal methods and capture complementary information from different data sources.
A prominent example is TwoStream-SLR [27], a dual-stream architecture designed for both signed language translation (SLT) and CSLR. It processes RGB and pose heatmaps in parallel using the first four blocks of S3D in each branch, with bidirectional lateral connections enabling multi-level information exchange. Additional supervision is provided through Sign Pyramid Networks and auxiliary prediction heads (video, keypoint, and joint). The model is trained using self-distillation, minimizing KL-divergence between the outputs of the three heads. An illustration of the architecture is shown in Fig. 16. This model has also been applied to ISLR [215], using a sliding window mechanism and leveraging the same framework to extract sign boundaries and construct training dictionaries.
Figure 16: Overview of the TwoStream-SLR [27]. Eight S3D blocks constitute the dual-stream multimodal architecture. Bidirectional Lateral Connection is utilized to exchange the information produced by each branch in a multi-level manner. The Sign Pyramid Networks, Video and Keypoint head are actins as auxiliary classifiers, providing extra supervision. Self-distillation is also employed to share the knowledge between the two streams.
Other CSLR models have adapted multimodal ISLR backbones. Uni-Sign [102] can be used directly for CSLR, while SignBERT+ [71] can be extended by appending temporal pooling and a CTC decoder to generate gloss sequences.
Beyond sign language, general-purpose CHGR systems have explored multimodality across broader visual sensor inputs. For example, [112] proposes a framework that segments RGB, depth, infrared, and optical flow inputs using a shared U-Net and processes them via separate 3D DenseNet-43 backbones. Extracted features are fused through Canonical Correlation Analysis to align shared representations across modalities.
6.4 State-of-the-art Methods for CSLR
This subsection reviews state-of-the-art (SOTA) methods for Continuous Sign Language Recognition (CSLR), focusing on their performance on the three most widely used benchmarks: PHOENIX14 [44], PHOENIX14T [26], and CSL-Daily [209].
As shown in Table 7, the dominant architecture remains the RGB-based Spatial Perception Model+1DCNN+BiLSTM pipeline, where the spatial perception model can be either a 2DCNN or a transformer-based approach. Typically, a ResNet18 backbone extracts spatial features, a 1D CNN or TCN models short-term temporal relations, and a BiLSTM captures long-range context for alignment. VAC-CSLR [122] established a widely adopted training setting to address optimization instability. Many works have since extended this baseline with auxiliary classifiers and additional loss terms to enhance feature learning [83, 4, 14]. Furthermore, the TwoStream-SLR [27] seems to compete worthily the aforementioned pipeline, surpassing it when extra training data are utilized [188] and ranked second in the CSL-Daily [209] benchmark without additional samples.
Recent innovations that improve CSLR include:
  • Data Augmentation Across Datasets: Leveraging visually similar signs from different corpora as training data [188].
  • Modeling Gloss Dependencies: Addressing the independence assumption in vanilla Connectionist Temporal Classification (CTC) loss through conditional modeling [197].
  • Multimodal Inputs: Combining RGB and pose heatmaps to improve robustness [27].
  • Graph-Based Feature Interaction: Using GNNs instead of CNNs to model intra- and inter-frame patch relationships [47].
  • Linguistic Priors: Integrating gloss embeddings or language models to guide learning [58, 57].
  • Pre-trained Vision Backbones: Fine-tuning foundation models like CLIP for richer feature extraction [14].
Table 7: Brief description of current state-of-the-art CSLR methods and their performance on the three most used benchmark (Phoenix 2014 [44], Phoenix 2014 T [26] and CSL-Daily [209]). Methods marked with star (*) have utilized additional training data.
Method Modality Architecture Description Training Setting Dataset WER
SLA-Adapter/LoRA [14] RGB CLIP visual encoder with adopter/LoRA, followed by TCN and BiLSTM VAC-CSLR [122] and early supervision, by adding an extra classifier after the visual encoder. CSL-Daily 25.8%
Phoenix 18.8%/19.3%
Phoenix-T 19.5%/19.4%
Uni-Sign [102]* RGB, Pose A pretrained LLM process features extracted by pose, visual and temporal encoders. Three-stage training, multi-task learning with additional training data CSL-Daily 26%
Swin-MSTP [15] RGB SwinT for spatial feature extraction and a mixture of 1D convolution, convolution with different dilation rates and TLP [73]. VAC-CSLR [122] CSL-Daily 27.1%
Phoenix 18.7%
Phoenix-T 19.7%
SlowFastSign [4] RGB Modified SlowFast with bidirectional feature fusion between the two streams. VAC-CSLR [122] and auxiliary classifiers to provide extra supervision. CSL-Daily 24.9%
Phoenix 18.3%
Phoenix-T 18.7%
MultiSignGraph [47] RGB Multiscale spatial-temporal GNN that considers image patches as the nodes. [47]. VAC-CSLR [122] CSL-Daily 26.4%
Phoenix 19.1%
Phoenix-T 19.1%
CSGC [143] RGB 2DCNN+1DCNN+BiLSTM VAC-CSLR [122], auxiliary losses to maintain inter- and cross-sentence consistency and gloss-prototype. CSL-Daily 26.7%
Phoenix 19%
Phoenix-T 19.5%
CorrNet+ [76] RGB Modification of the CorrNet [74], obtained by adding temporal attention after the correlation and identification modules. VAC-CSLR [122] and TLP [73]. CSL-Daily 28.2%
Phoenix 18.2%
Phoenix-T 19.1%
TCNet [111] RGB Combination of the TLP [73] and the CorrNet [74] VAC-CSLR [122] and an additional regularization term CSL-Daily 29.3%
Phoenix 18.9%
Phoenix-T 19.4%
C2ST [197] RGB SwinT+1DCNN+BiLSTM CGD loss, distillation to maintain the consistency between video- and local-level features and linguistic prior CSL-Daily 25.8%
Phoenix 17.7%
Phoenix-T 18.9%
CoSign-2s [83] Pose Dual stream approach based on ST-GCN, considering both joints and motions. Each stream consists of group-specific branches. Extra supervision via auxiliary classifier and complementary regularization to enhance the robustness to noise. CSL-Daily 27.2%
Phoenix 20.1%
Phoenix-T 20.1%
CorrNet [74] RGB 2DCNN equipped with the proposed identification and correlation modules, followed by 1DCNN+BiLSTM. VAC-CSLR [122] CSL-Daily 30.1%
Phoenix 19.4%
Phoenix-T 20.5%
[188]* RGB, Pose TwoStream-SLR [27] Utilization of extra training data via dictionary construction. Phoenix 16.7%
Phoenix-T 18.5%
TwoStream-SLR [27] RGB, Pose Dual stream (containing S3D blocks) with bidirectional lateral connections. Extra supervision achieved by auxiliary CTC losses and self-distillation applied to additional prediction heads. CSL-Daily 25.3%
Phoenix 18.8%
Phoenix-T 19.3%
7 Datasets
The performance and generalizability of hand gesture and sign language recognition (HGR/SLR) models are heavily influenced by the quality, diversity, and scale of the datasets used during training and evaluation. Over the years, a wide variety of datasets have been developed to support different HGR tasks, ranging from static and isolated gestures to continuous sign language recognition in real-world settings. These datasets vary widely in terms of the number of gestures or glosses, the types of modalities captured (e.g., RGB, depth, pose, IR), the presence of signer diversity, and the complexity of the sequences. In the following subsections, we categorize and review the most prominent datasets according to the specific recognition task they target: (i) Isolated and Continuous Sign Language Recognition, (ii) Semaphoric Gestures, and (iii) Other Gesture Recognition Tasks. This organization aims to provide a clear overview of the resources available for each task type and facilitate their appropriate use in future research.
7.1 Isolated and Continuous SLR
The datasets that have been developed to support video-based Sign Language Recognition (SLR) can be broadly categorized into those designed for Isolated Dynamic SLR (IDSLR) and Continuous SLR (CSLR). The former datasets typically focus on individual signs, recorded in controlled lab environments with a fixed vocabulary, making them well-suited for classification-based models. In contrast, continuous datasets present real-world sequences of signs, often extracted from broadcast material or spontaneous signing, requiring temporal segmentation and sequence modeling approaches. Additionally, some datasets support Sign Language Translation (SLT) by providing sentence-level gloss annotations and parallel spoken language transcripts.
The RWTH-PHOENIX-Weather 2014 [44] and its extended version PHOENIX-2014-T1 [26] are benchmark datasets offering sentence-level annotations in German Sign Language (DGS) captured from television weather forecasts. More recent efforts such as BOBSL [12] and OpenASL [160] have significantly expanded the signer base and vocabulary, with hours of British and American Sign Language footage collected from TV and web sources, respectively. Lab-recorded corpora like KETI [90], CSL-Daily [209], and How2Sign [41] incorporate multimodal inputs (e.g., pose, depth) and structured glosses, aiding both CSLR and SLT tasks. These datasets collectively enable the development and benchmarking of models under varying constraints of signer variability, language, modality, and task complexity.
On the other hand, isolated sign datasets such as MS-ASL [85], WLASL [99], and BSL-1K [11] provide extensive vocabularies (up to 2,000 signs) and diverse signers, facilitating the study of vocabulary scalability and signer generalization. Many of these datasets, such as AUTSL [164] and BosphorusSign22k [217], also include pose or depth modalities, enriching multimodal learning. While isolated datasets are primarily used for gesture classification, continuous datasets address the more complex problem of temporal segmentation, gloss prediction, and SLT. A few hybrid datasets like GSL SD/SI [2] and Corpus FinSL [154] provide data that can be used in both isolated and continuous settings. Together, these datasets serve a broad spectrum of research needs, from foundational gesture recognition to high-level translation across multiple sign languages.
Table 8: Overview of the characteristics of each dataset related to Sign Language Understanding (SLU) and Hand Gesture Recognition (HGR). The sign vocabulary (Sign Voc.) corresponds to the number of distinct classes included in each dataset.
Name Task Year Modality Samples Signers Sign Voc. Language Duration Source
PHOENIX-2014 [44] CSLR/SLT 2014 RGB - 9 1558 DGS 10.73h TV
PHOENIX-2014-T1 [26] CSLR/SLT 2018 RGB 8257 9 1066 DGS - TV
KETI [90] CSLR/SLT 2019 RGB, Pose 14672 14 524 KSL 27.99h Lab
GSL SD/SI [2] CSLR/SLT 2019 RGBD 10295 7 310 GSL 9.59h Lab
Corpus FinSL [154] CSLR/SLT 2020 RGB 343 21 - FinSL 14.37h -
CSL-Daily [209] CSLR/SLT 2021 RGB 20654 10 2000 CSL 23.27h Lab
How2Sign (green studio) [41] CSLR/SLT 2021 RGBD, Pose 2529 11 - ASL 79.12h Lab
BOBSL [12] CSLR/SLT 2021 RGB 1962 39 2281 BSL 1467h TV
OpenASL [160] CSLR/SLT 2022 RGB 98417 220 - ASL 288h Web
Auslan-Daily [158] CSLR/SLT 2023 RGB 25K 67 - AUSLAN - TV, Web
TVB-HKSL-News [130] CSLR/SLT 2024 RGB, Pose 7K 2 6515 HKSL 16.07h TV
QSL [48] CSLR/SLT 2024 RGB 30040 5 654 CSL - Lab
ChicagoFSWild+ [159] CSLR 2019 RGB 55232 260 26 ASL - Web
HKSL [212] CSLR 2021 RGBD 2400 6 55 HKSL - Lab
CSL [78] IDSLR 2018 RGBD, Pose 125K 50 500 CSL - Lab
SMILE [42] IDSLR 2018 RGB - 30 100 DSGS - Lab
ASL Dataset [16] IDSLR 2018 Pose 1200 20 30 ASL - Lab
MSASL [85] IDSLR 2019 RGB 25513 222 1000 ASL 24.65h Web
WLASL [99] IDSLR 2019 RGB 21083 119 2000 ASL 14.11h Web
GSL isol. [2] IDSLR 2019 RGBD 40785 7 310 GSL 6.44h Lab
BosphorusSign22k [217] IDSLR 2020 RGBD, Pose 22542 6 744 TSL 19h Lab
INCLUDE [169] IDSLR 2020 RGB 4287 7 263 ISL 3h Lab
AUTSL [164] IDSLR 2020 RGBD, Pose 38336 43 226 TSL 21h Lab
RKS-PERSIANSIGN [144] IDSLR 2020 RGB 10000 10 100 Persian - Lab
BSL-1K [11] IDSLR 2020 RGB 280K 40 1064 BSL 1060h TV
NMFs-CSL [69] IDSLR 2021 RGB 32010 10 1.067 CSL 24.4h Lab
KArSL [163] IDSLR 2021 RGBD, Pose 75300 3 502 ArSL - Lab
CSLD [184] IDSLR 2022 RGB 120K 30 400 CSL 66.8h Lab
NCSL [182] IDSLR 2022 RGB 90K 30 300 CSL 50.1h Lab
IISL2020 [92] IDSLR 2022 RGB 12100 16 11 ISL 7h Lab
LSA64 [149] IDSLR 2023 RGB 3200 10 64 LSA 1.9h Lab
Signsuisse Lexicon [126] IDSLR 2023 RGB 18221 - - 3 (Mult) - Web
FDMSE-ISL [135] IDSLR 2024 RGBD 40033 20 2002 ISL 36h Lab
7.2 Semaphoric HGR
Semaphoric hand gesture recognition (HGR) research is supported by a diverse range of datasets categorized into three main types based on task formulation: static, isolated dynamic, and continuous dynamic gesture recognition. These datasets vary in their application domains (e.g., Human-Computer Interaction (HCI), Human-Machine Interface (HMI), Human-Robot Interaction (HRI)), modalities (e.g., RGB, depth, pose, infrared), and in the number of gesture classes, subjects, and gesture samples they include. The statistics of datasets for semaphoric HGR are provided in tables 9,10 and 11, regarding the task for which they are used (i.e. static, isolated dynamic and continuous HGR, respectively).
Static HGR datasets focus on recognizing gestures from single-frame input and are generally easier to process due to the absence of temporal complexity. Notable examples include LaRED [67], one of the earliest RGBD datasets with 243,000 samples and 81 classes, and HaGRIDv2 [131], a large-scale RGB dataset designed for HMI with over 1 million samples, covering 34 classes. Smaller yet targeted datasets such as OUHANDS [118] and URGR [17] focus on specific domains like HCI and HRI, while MD-UHGRD [194] offers moderately sized samples and class diversity. Isolated dynamic HGR datasets contain gestures performed in isolation, with start and end points explicitly segmented. Datasets like ChaLearn IsoGD [179] and EgoGesture (isolated) [199] feature rich modality (RGBD, IR, pose) and high gesture diversity, with up to 249 classes. Others like Jester [117] offer large-scale RGB-only data from over 1,300 subjects, making them suitable for data-driven deep learning approaches. More recent contributions such as DATE [35] and ZJUGesture [193] reflect growing interest in real-world HCI scenarios. Continuous dynamic HGR datasets challenge models to identify and segment gestures from continuous video streams. The ChaLearn ConGD [179] dataset remains a benchmark with 22,535 sequences and 249 gesture classes. Similarly, EgoGesture [199] and IPN Hand [19] include both isolated and continuous segments, enabling comprehensive model development and evaluation. Datasets like LD-ConGR [106] and ZJUGesture [193] contribute additional complexity by increasing subject diversity and contextual variability.
Each dataset supports specific sub-tasks within HGR research: static datasets aid in gesture classification, isolated dynamic datasets support temporal gesture modeling and classification, and continuous datasets allow for joint gesture spotting and classification—crucial for real-time and interactive applications. This diversity in dataset structure, scale, and modality reflects the evolving demands and challenges of semaphoric gesture recognition across interactive domains.
Table 9: Datasets for static hand gesture recognition.
Name Year Domain Modality Samples Subjects Classes Resolution
HGR-1 [88] 2014 - RGB 899 12 27 174x131-640x480
LaRED [67] 2014 - RGBD 243000 10 81 640x480
OUHANDS [118] 2016 HCI RGBD 3150 23 10 640x480
LRHG [211] 2021 HRI RGB 4320 8 10 640x480
SHAPE [33] 2022 - RGB 34226 20 32 4128x3096
HaGRID [86] 2024 HCI RGB 552992 34730 19 1920x1080
HaGRIDv2 [131] 2024 HCI RGB 1086158 65977 34 1920x1080
MD-UHGRD [194] 2024 HRI RGB 20000 - 19 -
URGR [17] 2024 HRI RGB 347483 16 6 480x640
Table 10: Datasets for isolated dynamic hand gesture recognition.
Name Year Domain Modality Samples Subjects Classes Frames
SKIG [108] 2013 - RGBD 1080 6 10 -
nvGesture [124] 2016 HMI RGBD, IR, OF 1532 20 25 -
ChaLearn IsoGD [179] 2016 - RGBD 47933 21 249 -
DHG-14/28 [37] 2016 - Depth, Pose 2800 20 14/28 -
SHREC17-14/28 [166] 2017 - Depth, Pose 2800 28 14/28 -
EgoGesture (isol.) [199] 2017 HCI RGBD 24161 50 83 2953224
Briareo [115] 2019 HMI RGBD, Pose, IR 1440 40 12 -
Jester [117] 2019 HCI RGB 148092 1376 27 5331312
IPN Hand (isol.) [19] 2020 HCI RGB 4218 50 13 -
TCG [189] 2020 HMI Pose 250 5 4 839350
LD-ConGR (isol.) [106] 2022 HCI RGBD 44887 30 10 -
ZJUGesture (isol.) [193] 2023 HCI RGB 9892 60 9 -
DATE [35] 2024 HCI RGB 13500 22 27 -
EgoDriving [138] 2024 HMI RGB 400 - 29 70276
Table 11: Datasets for continuous dynamic hand gesture recognition.
Name Year Domain Modality Samples Instances Subjects Classes
ChaLearn ConGD [179] 2016 - RGBD 22535 47933 21 249
EgoGesture (cont.) [199] 2017 HCI RGBD 2081 24161 50 83
IPN Hand (cont.) [19] 2020 HCI RGB 200 4218 50 13
Police Gesture Dataset [66] 2020 HMI RGB 21 3354 - 8
LD-ConGR (cont.) [106] 2022 HCI RGBD 542 44887 30 10
ZJUGesture (cont.) [193] 2023 HCI RGB - 9892 60 9
7.3 Other
In addition to standard semaphoric gesture datasets, several specialized datasets have been introduced to support research on more specific gesture types. The First-Person Hand Action (FPHA) dataset [50], published in 2018, contains 1,175 video sequences captured from a first-person perspective by six participants. It focuses on isolated manipulative hand gestures across 45 action classes, offering RGB, depth, and skeleton data, and is particularly valuable for studies involving egocentric interaction. Another notable example is the DP Dataset [127], released in 2023, which, to the best of our knowledge, is the only publicly available benchmark specifically designed for pointing gesture recognition. Collected from 33 participants in two indoor environments, the dataset comprises approximately 2.8 million frames and 6,335 annotated gesture instances. It includes detailed annotations such as gesture timings and pointing directions, enabling both binary gesture classification (gesture vs. non-gesture) and directional estimation tasks.
8 Evaluation Metrics
In this section, we present the key evaluation metrics adopted in the literature for measuring the performance of hand gesture recognition (HGR) methods. Broadly, metrics fall into two main categories: classification metrics (used in static and isolated dynamic HGR) and sequence-to-sequence metrics (used in continuous dynamic HGR). Additionally, other indicators such as computational resource usage and overfitting-related metrics are also considered.
8.1 Classification Metrics
In multi-class HGR problems—especially when dealing with class imbalance—researchers often report both micro-averaged (per-instance) and macro-averaged (per-class) versions of common classification metrics. Micro-averaged metrics aggregate contributions from all instances across classes, thus being sensitive to class imbalance. Macro-averaged metrics compute the metric independently for each class and then average, treating all classes equally regardless of their frequency.
Micro-averaged metrics:
$\text{Accuracy}_{micro}=\frac{\sum_{i=1}^{N}(TP_{i}+TN_{i})}{\sum_{i=1}^{N}(TP_{i}+TN_{i}+FP_{i}+FN_{i})}$ (1)
$\text{Precision}_{micro}=\frac{\sum_{i=1}^{N}TP_{i}}{\sum_{i=1}^{N}(TP_{i}+FP_{i})}$ (2)
$\text{Recall}_{micro}=\frac{\sum_{i=1}^{N}TP_{i}}{\sum_{i=1}^{N}(TP_{i}+FN_{i})}$ (3)
[…]
Macro-averaged metrics:
$\text{Accuracy}_{macro}=\frac{1}{N}\sum_{i=1}^{N}\frac{TP_{i}+TN_{i}}{TP_{i}+FP_{i}+TN_{i}+FN_{i}}$ (5)
$\text{Precision}_{macro}=\frac{1}{N}\sum_{i=1}^{N}\frac{TP_{i}}{TP_{i}+FP_{i}}$ (6)
$\text{Recall}_{macro}=\frac{1}{N}\sum_{i=1}^{N}\frac{TP_{i}}{TP_{i}+FN_{i}}$ (7)
[…]
F1-Score: For both micro and macro cases, the F1-score is defined as the harmonic mean of Precision and Recall:
$F_{1}=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ (9)
Here, $TP_{i}$, $FP_{i}$, $TN_{i}$, and $FN_{i}$ denote the number of true positives, false positives, true negatives, and false negatives for the $i$-th class, respectively, and $N$ is the total number of classes.
8.2 Sequence-to-Sequence Metrics
For continuous HGR (CHGR), particularly in Continuous Sign Language Recognition (CSLR), sequence-level metrics are used. Word Error Rate (WER) [122, 2] is the standard metric, measuring the minimum number of substitutions ($S$), deletions ($D$), and insertions ($I$) needed to transform the predicted sequence into the ground truth sequence of length $N$:
$WER=\frac{S+D+I}{N}$ (10)
Edit Distance Accuracy [45, 185] is a complementary metric representing the ratio of correctly predicted elements in the sequence:
[…]
These metrics provide insight not only into classification accuracy but also into temporal and sequential consistency.
8.3 Computational Resource Metrics
In addition to recognition performance, computational efficiency is often evaluated, especially for real-time or embedded HGR applications. Common metrics include:
  • Inference Time: Time (in ms or seconds) taken to process one input sample.
  • Frames Per Second (FPS): Number of frames processed per second; higher values indicate faster processing.
  • Model Size: The number of trainable parameters or total memory footprint of the model.
These metrics are crucial when deploying HGR systems on edge devices or in latency-sensitive environments.
8.4 Overfitting Indicators
To assess generalization and detect overfitting, the following indicators are used:
  • Train-Test Performance Gap: A significant difference between training and testing accuracy (or loss) often signals overfitting.
  • Validation Loss Trend: Increasing validation loss while training loss decreases is a common overfitting pattern.
  • Early Stopping Criteria: Monitored via a plateau or increase in validation loss across epochs.
9 Current Challenges and Future Research Directions
As mentioned in the Methods sections (Sections 4, 5, and 6), researchers have developed sophisticated techniques and large-scale datasets to achieve satisfactory performance. However, as shown in the relevant tables (Tables 3, 4, 5, 6, 7), recognition accuracy remains low in many cases—particularly in more complex tasks such as Continuous Sign Language Recognition (CSLR). In this section, we discuss the current challenges and limitations that hinder efficient Human Gesture Recognition (HGR), posing ongoing concerns for the research community and shaping future research directions. Additionally, we review the techniques proposed to address these issues, offering insights that may guide future efforts to overcome the identified challenges.
9.1 Gesture-Irrelevant Factors
The presence of gesture-irrelevant factors (or redundant information) hinders accurate recognition in diverse real-world scenarios. Some of these factors include the following [171, 9, 80, 17, 141, 206]:
  • Environmental factors: Cluttered background, varying illumination conditions, and self-occlusions make it difficult or even impossible to automatically spot the hand region, which is an important step for efficient HGR;
  • Subject characteristics: Subject independence –—i.e., the ability to correctly identify a gesture regardless of the signer’s characteristics such as skin color, hand shape variations, or clothing—is a key prerequisite for achieving sufficient generalization and enabling real-world deployment.
  • Variations in perception device position: The distance between the performer and the camera angle may reduce the recognition performance;
  • Input data quality: Low resolution and blurriness may corrupt the discriminative characteristics of the hand gesture region.
Aiming to address these challenges, researchers have proposed segmentation models to crop hand regions and remove background interference [208, 34, 153], object detectors for spatial normalization [17, 9], super-resolution techniques to enhance image quality [17], attention mechanisms [206, 207] to enforce the visual encoder to pay attention on the most informative regions, information-theoretic losses [100], and feature selection techniques [173] to reduce information redundancy. Many of these approaches rely on computationally expensive off-the-shelf models, which increase the complexity of the overall HGR methods and set an upper bound to their accuracy. Thus, the presence of gesture-irrelevant factors could concern future research work, while potentially being combined with eXplainable Artificial Intelligence (XAI) approaches [148] to facilitate this task.
9.2 Data Scarcity
Many existing datasets either contain a relatively small number of samples or lack sufficient diversity. For example, among static HGR datasets, only the versions of HaGRID [86, 131] have been collected from more than one hundred participants, despite some others having a plausible number of samples [67, 17]. A similar situation is observed in IDHGR datasets, where only the Jester dataset [117] includes data from over one hundred subjects. Furthermore, the amount of samples for specific modalities, such as IR data, is limited compared with the size of RGB-based datasets, an issue known as data imbalance [181]. The lack of large-scale, diverse datasets leads to overfitting and limited generalization, which is closely related to the challenges discussed earlier. This issue is further exacerbated in tasks that involve costly annotation processes and require data-intensive deep learning models with a large number of parameters, such as CSLR.
Several techniques have been explored to mitigate this challenge:
  • Multilingual Data Aggregation: Combining datasets (e.g., CSL and DGS) to exploit cross-lingual visual similarities increases the overall amount of training data [188].
  • Self-Supervised Learning: In the context of the self-supervised learning setting [91], gloss-to-gloss translation tasks have been used for parameter initialization [205], providing linguistic priors to sign language models. Pretraining with pose reconstruction has also been investigated to incorporate prior knowledge of hand structure [71, 68, 79].
  • Advanced Data Augmentation: Adversarial learning has been employed to optimize augmentation hyperparameters [128].
  • Few-/Zero-shot Learning: Approaches using only a small number of samples to correctly recognize gestures—often based on CLIP—have also been proposed [216, 24, 147, 146, 22].
These approaches have led to improved recognition performance, even when trained on limited data. However, accuracy remains suboptimal, making data scarcity a persistent research challenge and a key direction for future work in the field.
9.3 Optimization Difficulties
Training the visual encoder effectively remains difficult, especially in deep architectures for CSLR. As shown in CTCA [57] (Fig. 17), increasing the depth of the temporal module (e.g., 1D-TCN followed by BiLSTM) improves its generalization but hampers the encoder’s learning due to gradient attenuation in backpropagation. Furthermore, the CTC spike phenomenon—where the model over-focuses on a few keyframes—can cause rapid overfitting and limit encoder training [122, 62].
Figure 17: Experimental results of a study conducted by authors in [57], which aims to measure the impact of the depth of the temporal aggregation module (a 1D-TCN with various number of layers followed by a two-layer BiLSTM) on the effectiveness of training the visual encoder (aka spatial perception module). In both diagrams, the x axis represents the number of layers of the 1D-TCN. In the top chart, the two modules are compared considering the information stored in weights (IIW) [187] (y axis), which can be viewed as an indicator of generalization gap, while in the bottom the corresponding WER is shown. As the number of layers of the 1D-TCN increases, although the generalization ability of the short-term temporal module increases, the overall performance decreases due to the insufficient training of the visual encoder.
To address these issues, several strategies have been proposed:
  • Auxiliary Supervision: Adding early-stage classifiers to the encoder or short-term module has been proposed as a strategy to guide training by providing additional supervision [122, 143, 74, 76, 111, 213].
  • Temporal Module Redesign: Placing the 1D-TCN and BiLSTM in parallel instead of sequentially [57].
  • Linguistic Constraints and Gloss Segmentation: Using gloss boundaries [62] or prior knowledge [58] to regularize learning.
The aforementioned approaches —particularly auxiliary supervision [122]— have served as the foundation for many existing methods. However, the increasing complexity of modern deep architectures often exceeds the effectiveness of these techniques. As a result, optimization remains a significant challenge —one that may be further exacerbated by the continued development of even deeper models.
9.4 Limitations of Architectures
Despite the achievements of DL, the existing architectures suffer from the following limitations:
  • Block-fixed models: The existing architectures are based on hand-crafted design, which may not be optimal[206].
  • GCNs with fixed topology graphs: Standard GCNs use fixed topology graphs based on hand anatomy, which may not capture optimal inter-joint relationships.
  • Ineffective multimodal fusion: Most of the current works utilize late fusion, that does not fully leverage the available information [68, 71, 202, 204, 203, 104, 82, 81].
To address these limitations, researchers have proposed various strategies, including Neural Architecture Search (NAS) to identify optimal model structures [206], learnable or adaptive skeleton adjacency matrices [77, 30, 28, 107], and multi-level fusion for dual-stream approaches [60]. Despite these advances, several aspects of deep learning architectures remain underexplored, leaving a substantial research gap.
10 Conclusions
The undeniable importance of vision-based hand gesture recognition (HGR) stems from its wide range of applications—spanning human-robot interaction to sign language understanding. This broad applicability has made it a vibrant and active research field. Recent advances in deep learning have played a pivotal role in driving progress, enabling the development of increasingly sophisticated architectures and training strategies aimed at tackling more complex and fine-grained tasks, such as continuous sign language recognition.
This survey focused on HGR methods that leverage visual input and deep learning. These methods were highlighted not only for their superior performance compared to traditional hand-crafted approaches, but also for their higher potential for deployment in real-world scenarios, unlike more cumbersome sensor-based techniques. From the surveyed literature, it becomes clear that even relatively simple and shallow deep learning models can achieve high accuracy in static HGR tasks. In such cases, the primary challenges are typical of general computer vision problems and are not inherently tied to HGR itself.
Despite the significant progress achieved in vision-based hand gesture recognition, several open challenges remain that shape future research directions. A major concern is the generalization of models across diverse environments, lighting conditions, camera viewpoints, and user-specific variations, which often lead to performance drops in real-world settings. The lack of large-scale, diverse, and well-annotated datasets, especially for continuous gestures and sign language, also hinders the development of more robust models. Furthermore, model efficiency and scalability are critical for deployment on resource-constrained edge devices, necessitating advances in lightweight architectures and model compression techniques. The integration of multimodal data, such as combining vision with audio or inertial signals, remains underexplored yet promising. Finally, interpretability and trustworthiness of deep learning models are essential, particularly for safety-critical applications like healthcare or assistive technologies. Addressing these challenges will not only enhance the practical applicability of HGR systems but also contribute to more inclusive and accessible human-computer interaction paradigms.
[…]
Acknowledgments The research leading to these results received funding from the European Commission under Grant Agreement No. 101168042 (TRIFFID) and No. 101189557 (TORNADO).
Declarations
Funding The research leading to these results received funding from the European Commission under Grant Agreement No. 101168042 (TRIFFID) and No. 101189557 (TORNADO).
Conflict of interest The authors have no competing interests to declare that are relevant to the content of this article.
Author contribution Konstantinos Foteinos, Jorgen Cani, Manousos Linardakis, Iraklis Varlamis and Georgios Th. Papadopoulos performed the literature review and prepared a draft of the manuscript. Panagiotis Radoglou-Grammatikis, Vasileios Argyriou and Panagiotis Sarigiannidis performed reviewing and editing of the manuscript. Georgios Th. Papadopoulos was responsible for receiving the funding to implement the study.